# AdaAugment: A Tuning-Free and Adaptive Approach to Enhance Data Augmentation

### v2

@@ 5 Conclusion
In this work, we propose AdaAugment, an innovative and tuning-free method for adaptive data augmentation. By adaptively adjusting the magnitudes of DA operations based on the real-time feedback obtained during the training of the target model, AdaAugment optimizes the data variability. This dynamic adjustment ultimately mitigates the risks associated with both underfitting and overfitting, thereby optimizing model training and enhancing model generalization capabilities. Through extensive empirical evaluations, we demonstrate that our adaptive augmentation mechanism helps to significantly improve the performance across various benchmark datasets with competitive efficiency. In the future, we hope to generalize AdaAugment to more tasks.
6 Impact Statements
The method proposed in this paper, grounded in both theory and experimentation, aims to advance research in data augmentation within the field of deep learning. The experimental data is sourced entirely from standard datasets, eliminating any ethical concerns. Additionally, there are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.

### v3

@@ VI Conclusion
In this work, we propose AdaAugment, an innovative and tuning-free method for adaptive data augmentation. By adaptively adjusting the magnitudes of DA operations based on the real-time feedback obtained during the training of the target model, AdaAugment optimizes the data variability. This dynamic adjustment ultimately mitigates the risks associated with both underfitting and overfitting, thereby optimizing model training and enhancing model generalization capabilities. Through extensive empirical evaluations, we demonstrate that our adaptive augmentation mechanism helps significantly improve performance across various benchmark datasets with competitive efficiency.
Acknowledgments
This work is supported by the STI 2030-Major Projects of China under Grant 2021ZD0201300, the Fundamental Research Funds for the Central Universities under Grant 2024300394, the National Natural Science Foundation of China under Grant 62276127.